Ry(X) denotes X roundedto the nearest y's place.what is the value of R10 (3,619)+R100(3,619)+R1.000(9,619)
Answer by Earlsdon(6103)

R10(3,619) = 3,620
R100(3,619) = 3,600
R1,000(9,619) = 10,000
Then: R10(3,619)+R100(3,619)+R1,000(9,619) = 3,620+3,600+10,000 = 17,220